OpenStudy (anonymous):

Cormac is making 27 phone calls each week as part of a fundraiser. Last week, he made 12 phone calls on weekdays. Which equation could be used to find x, the number of phone calls he made last week on weekend days?

OpenStudy (bahrom7893):

27-12=x

OpenStudy (anonymous):

27 = 12+x x = weekend calls